# Complexity of Quadratic Bosonic Hamiltonian Simulation:

### 27 Mar 2026 v1

@@ Abstract
where $A,C\in\operatorname{Mat}(M\times M,\mathds{R})$ are symmetric $d$-sparse matrices, i.e. each row and column contains at most $d$ nonzero entries. For now, we require $A$ and $C$ to be positive semidefinite (PSD), amounting to a lower semibounded spectrum of the bosonic Hamiltonian. Further details, including extensions beyond the PSD case, are provided in Appendix C.2.
For our construction, we impose a sign convention on the elements of $A$, motivated by a potential energy arising from spring-like coupling, i.e., of the form
@@
with coupling constants $l_{jk}\geq 0$ for all $j,k$. This is equivalent to requiring that $A$ and $C$ are Laplacian-type stiffness matrices, which are positive semidefinite and diagonally dominant. Note that this specific sign convention (beyond $A$ and $C$ being PSD) is not required for $\mathsf{BQP}$-completeness, but simplifies the construction of the simulation algorithm.
@@
Starting from an oracle for the $d$-sparse coupling matrices $A$ and $C$, one can build $d$-sparse oracles for $B,B^{\dagger}$ and $D,D^{\dagger}$ [8]. With this, it is straightforward to construct an oracle for the $rd$-sparse Hamiltonian on the right hand side of Eq. (12). Evolution under sparse Hamiltonians can be efficiently simulated, for instance using one of the methods in [9, 10, 11, 12]. In direct analogy to [2], we can ensure a final error $\varepsilon$ using $Q=\mathcal{O}(tKd\left\lVert H_{0}\right\rVert_{\rm max}+\log(1/\varepsilon))$ queries of $H_{0}$ and a number of gates $G=\mathcal{O}(Q\log^{2}(\frac{MQ}{\varepsilon}))$.
@@
over $(\alpha_{1},...,\alpha_{r})\in I_{0}$ for an efficiently characterizable index set $I_{0}$. The normalization factor $N$ from Eq. (10) is set by the initial conditions and remains unchanged under unitary time evolution.
@@
Note that $\tilde{T}$, and thus $A$ and $C$, have precisely the structure required for a network of harmonically coupled oscillators: Its off-diagonal elements are non-positive,
@@
Other families of Hamiltonians with $E^{+}=0$ can also implement postselection. As we show in Appendix B.7, negative eigenvalues in $A$ or mixed terms from $E^{-}\neq 0$ can be used to create the same dynamics as in the aforementioned family with $E^{+}\neq 0$.
@@ Appendix A Post-Processing of Measurement Data
In practice, one can gain efficiency when estimating all relevant amplitudes by recycling measurement outcomes. As measurements may also yield outcomes corresponding to coordinates $z_{(a,b)}$, which encode relative positions and momenta, these additional samples can be incorporated to reduce estimator variance. We restrict the discussion to the cases $r=1$ and $r=2$ and to positional coordinates $q$, governed by the matrix $A$. The analogous construction for momenta follows by replacing $q\to p$, $A\to C$, and $B\to D$.
@@
The prefactor $\sqrt{A_{jj}+\sum_{k\neq j}A_{jk}}$ can be computed using a single row query of $A$. Measuring the probability of the computational basis state $\left|(j,j)\right\rangle$ therefore allows to compute
@@ B.1 PostBQP-completeness Construction
over an efficiently characterizable index set $I$, up to $1/\mathrm{polylog}(M)=1/\mathrm{poly}(n)$ precision. For our reduction, we need to perform two such read-outs after a suitably chosen time $t_{f}=O(L)$. They are given by the index sets
@@ B.3 Separation of Quantum Circuit Logic and Dynamics on the Clock Register
Importantly, note that due to the way in which $S$ is constructed, the state of the computational registers $PQR$ at time $l=L+1$, which is what we choose for the read-out, is the same in the original and the transformed basis, such that we can analyze the read-out in the new basis.
@@ B.7 Alternative PostBQP-hard families of bosonic Hamiltonians
with $A_{FK}$ being the matrix defined in Eq. (28) and $Z$ and $Y$ are Pauli matrices. One can directly check that (1.) $E^{-}$ and $C$ anticommute, (2.) $CA$ describes two copies of the same Feynman-Kitaev Hamiltonian and (3.) $(E^{-})^{2}=-\beta^{2}\Pi_{L+2,1_{P}}\otimes\mathds{1}$ postselects onto the last readout layer in both copies.
@@
reproduces the generator used in the proof above. It has the same spectral properties – as before – and when starting in the analog initial state with only $\left\langle p^{(l=1)}_{0}(0)\right\rangle=1$, one can solve the postselected circuit problem by estimation of the momenta in the output layer. Note that while $A$ and $C$ individually are indefinite, the product $CA$ only has positive eigenvalues, which shows that the non-unitary dynamics necessary for postselection comes solely from $(E^{-})^{2}$, cf. also appendix C.2.
A third family of Hamiltonians which are $\mathsf{PostBQP}$-hard is given by $E=0$, but lifting the condition that $CA$ has to have a positive spectrum, as we require in systems of inertially coupled bosons. In this family of Hamiltonians one can directly choose $A$ to be the targeted generator $\tilde{A}$ as in Eq. (33) with a negative eigenvalue, which is now allowed.
@@ C.1 Most General Quadratic Bosonic Hamiltonian
C.2 Positivity of coupling matrices $A$ and $C$
As we show in the main text, coordinates can be chosen as $B^{\dagger}q$ and $D^{\dagger}p$ for $A=BB^{\dagger}$ and $C=DD^{\dagger}$, such that correlations evolve unitarily in the absence of mixed terms, $E=0$. If in addition to positivity of $A$ and $C$, the matrix elements are derived from positive coupling constants (cf. Eqs. (17) and (18)), an explicit and efficiently queriable form of $B$ and $D$ is known, cf. Eq. (7). If this sign convention is not met, an oracle access to quantum circuits implementing $\sqrt{A}$ and $\sqrt{C}$ can be employed to simulate time evolution via quantum phase estimation, cf. Theorem 4 of [2].
In fact, the condition that both, $A$ and $C$, are PSD is stronger than necessary. In the case $E=0$, a positive spectrum of the product $AC$ is sufficient to ensure unitary evolution, cf. Eq. (11). This can be achieved using indefinite or even negative $A$ and $C$ if they share the same block structure in a decomposition into positive and negative part such that signs are correctly canceled in the product. An example that deviates from this condition is the single-mode squeezing Hamiltonian

### 14 Apr 2026 v2

@@ Abstract
where $A,C\in\operatorname{Mat}(M\times M,\mathds{R})$ are symmetric $d$-sparse matrices, i.e. each row and column contains at most $d$ nonzero entries. For now, we require AA and $C$ to be positive semidefinite (PSD), amounting to a lower semibounded spectrum of the bosonic Hamiltonian. Further details, including extensions beyond the PSD case, are provided in Appendix C.2.
For our construction, we impose a sign convention on the elements of AA, motivated by a potential energy arising from spring-like coupling, i.e., of the form
@@
with coupling constants $l_{jk}\geq 0$ for all $j,k$. This is equivalent to requiring that AA and $C$ are Laplacian-type stiffness matrices, which are positive semidefinite and diagonally dominant. Note that this specific sign convention (beyond AA and $C$ being PSD) is not required for $\mathsf{BQP}$-completeness, but simplifies the construction of the simulation algorithm.
@@
Starting from an oracle for the $d$-sparse coupling matrices AA and $C$, one can build $d$-sparse oracles for $B,B^{\dagger}$ and $D,D^{\dagger}$ [8]. With this, it is straightforward to construct an oracle for the rdrd-sparse Hamiltonian on the right hand side of Eq. (12). Evolution under sparse Hamiltonians can be efficiently simulated, for instance using one of the methods in [9, 10, 11, 12]. In direct analogy to [2], we can ensure a final error $\varepsilon$ using $Q=\mathcal{O}(tKd\left\lVert H_{0}\right\rVert_{\rm max}+\log(1/\varepsilon))$ queries of $H_{0}$ and a number of gates $G=\mathcal{O}(Q\log^{2}(\frac{MQ}{\varepsilon}))$.
@@
over $(\alpha_{1},...,\alpha_{r})\in I_{0}$ for an efficiently characterizable index set $I_{0}$. The normalization factor NN from Eq. (10) is set by the initial conditions and remains unchanged under unitary time evolution.
@@
Note that $\tilde{T}$, and thus AA and $C$, have precisely the structure required for a network of harmonically coupled oscillators: Its off-diagonal elements are non-positive,
@@
Other families of Hamiltonians with $E^{+}=0$ can also implement postselection. As we show in Appendix B.7, negative eigenvalues in AA or mixed terms from $E^{-}\neq 0$ can be used to create the same dynamics as in the aforementioned family with $E^{+}\neq 0$.
@@ Appendix A Post-Processing of Measurement Data
In practice, one can gain efficiency when estimating all relevant amplitudes by recycling measurement outcomes. As measurements may also yield outcomes corresponding to coordinates $z_{(a,b)}$, which encode relative positions and momenta, these additional samples can be incorporated to reduce estimator variance. We restrict the discussion to the cases $r=1$ and $r=2$ and to positional coordinates $q$, governed by the matrix AA. The analogous construction for momenta follows by replacing $q\to p$, $A\to C$, and $B\to D$.
@@
The prefactor $\sqrt{A_{jj}+\sum_{k\neq j}A_{jk}}$ can be computed using a single row query of AA. Measuring the probability of the computational basis state $\left|(j,j)\right\rangle$ therefore allows to compute
@@ B.1 PostBQP-completeness Construction
over an efficiently characterizable index set II, up to $1/\mathrm{polylog}(M)=1/\mathrm{poly}(n)$ precision. For our reduction, we need to perform two such read-outs after a suitably chosen time $t_{f}=O(L)$. They are given by the index sets
@@ B.3 Separation of Quantum Circuit Logic and Dynamics on the Clock Register
Importantly, note that due to the way in which $S$ is constructed, the state of the computational registers PQRPQR at time $l=L+1$, which is what we choose for the read-out, is the same in the original and the transformed basis, such that we can analyze the read-out in the new basis.
@@ B.7 Alternative PostBQP-hard families of bosonic Hamiltonians
with $A_{FK}$ being the matrix defined in Eq. (28) and $Z$ and $Y$ are Pauli matrices. One can directly check that (1.) $E^{-}$ and $C$ anticommute, (2.) CACA describes two copies of the same Feynman-Kitaev Hamiltonian and (3.) $(E^{-})^{2}=-\beta^{2}\Pi_{L+2,1_{P}}\otimes\mathds{1}$ postselects onto the last readout layer in both copies.
@@
reproduces the generator used in the proof above. It has the same spectral properties – as before – and when starting in the analog initial state with only $\left\langle p^{(l=1)}_{0}(0)\right\rangle=1$, one can solve the postselected circuit problem by estimation of the momenta in the output layer. Note that while AA and $C$ individually are indefinite, the product CACA only has positive eigenvalues, which shows that the non-unitary dynamics necessary for postselection comes solely from $(E^{-})^{2}$, cf. also appendix C.2.
A third family of Hamiltonians which are $\mathsf{PostBQP}$-hard is given by $E=0$, but lifting the condition that CACA has to have a positive spectrum, as we require in systems of inertially coupled bosons. In this family of Hamiltonians one can directly choose AA to be the targeted generator $\tilde{A}$ as in Eq. (33) with a negative eigenvalue, which is now allowed.
@@ C.1 Most General Quadratic Bosonic Hamiltonian
C.2 Positivity of coupling matrices AA and $C$
As we show in the main text, coordinates can be chosen as $B^{\dagger}q$ and $D^{\dagger}p$ for $A=BB^{\dagger}$ and $C=DD^{\dagger}$, such that correlations evolve unitarily in the absence of mixed terms, $E=0$. If in addition to positivity of AA and $C$, the matrix elements are derived from positive coupling constants (cf. Eqs. (17) and (18)), an explicit and efficiently queriable form of $B$ and $D$ is known, cf. Eq. (7). If this sign convention is not met, an oracle access to quantum circuits implementing $\sqrt{A}$ and $\sqrt{C}$ can be employed to simulate time evolution via quantum phase estimation, cf. Theorem 4 of [2].
In fact, the condition that both, AA and $C$, are PSD is stronger than necessary. In the case $E=0$, a positive spectrum of the product ACAC is sufficient to ensure unitary evolution, cf. Eq. (11). This can be achieved using indefinite or even negative AA and $C$ if they share the same block structure in a decomposition into positive and negative part such that signs are correctly canceled in the product. An example that deviates from this condition is the single-mode squeezing Hamiltonian